# Strong solutions for a class of stochastic thermo-magneto-hydrodynamic-type systems with multiplicative noise

### v1

@@ 3.3. Stochastic Boussinesq model for the Bénard convection
Since $\mathcal{S}$ has the same form as (3.4), it remains to verify (2.7), (2.8), (2.9), and (2.10) for the map $\mathcal{F}$. By Hölder’s inequality and Sobolev embedding,
@@
which implies (2.7). Inequality (2.8) can be shown similarly. Next, we will prove (2.9) and (2.10) for the map $\mathcal{F}$. By Hölder’s inequality and Sobolev embedding,
$\left|{\mathcal{F}(\Phi)}\right|^{2}$ $\leq C\left|{\theta}\right|^{2}+C\left|{u}\right|_{L^{4}}^{2}\left|{\nabla\phi}\right|_{L^{4}}^{2}\leq C\left\|{\Phi}\right\|^{2}.$
By interpolation and Gagliardo–Nirenberg inequalities,
@@
$\leq C\left\|{\Phi}\right\|\left|{\mathcal{A}\Phi}\right|.$
These inequalities imply (2.9) and (2.10), thus completing the verification of condition (R).
@@ 3.5. Stochastic micropolar and magneto-micropolar fluid systems
The existence of strong solutions to the determinisitic magneto-micropolar fluid system is shown in [40, 49]. In the stochastic case, the existence of martingale weak solutions to the system is proven in [50, 51]. This paper establishes the existence and uniqueness of local pathwise strong solutions to the stochastic magneto-micropolar fluid system for $d\leq 3$. This pathwise strong solution is shown to be global when $d\leq 2$.
@@ 4. Existence and uniqueness of a local strong solution
Let $\mathcal{S}$ be a given stochastic basis. The pair $(\Phi,\tau)$ is a local weak pathwise solution of (1.1) with initial data $\Phi_{0}$ if $\tau$ is a strictly positive stopping time and $\Phi(\,\cdot\wedge\tau)$ is a predictable process in $\mathcal{V}^{\prime}$ such that
@@
$\Phi(t\wedge\tau)+\int_{0}^{t\wedge\tau}\big(\mathcal{A}\Phi+\mathcal{B}(\Phi)+\mathcal{R}\Phi\big)\,\mathrm{d}s=\Phi_{0}+\int_{0}^{t\wedge\tau}g(\Phi)\,\mathrm{d}W\quad\text{in $\mathcal{V}^{\prime}$}.$ (4.2)
@@
Let $\mathcal{S}$ be a given stochastic basis. A local weak pathwise solution $(\Phi,\tau)$ is said to be a local strong pathwise solution of (1.1) with initial data $\Phi_{0}$ if $\Phi(\,\cdot\wedge\tau)$ is a predictable process in $\mathcal{H}$ such that
@@
and that for any $t>0$, the equality (4.2) is satisfied in $\mathcal{H}$.
@@
$\leq C\lambda_{n}^{-1}\left\|{Q_{n}\mathcal{B}(\Phi^{n})}\right\|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-1}\left\|{\Phi^{n}}\right\|\left|{\mathcal{A}\Phi^{n}}\right|^{3}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-\frac{1}{2}}\left\|{\Phi^{n}}\right\|^{2}\left|{\mathcal{A}\Phi^{n}}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}.$ (4.13)
@@
$\leq C\lambda_{n}^{-1}\left\|{Q_{n}\mathcal{R}(\Phi^{n})}\right\|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-1}\left(1+\left\|{\Phi^{n}}\right\|^{4}\right)\left|{\mathcal{A}\Phi^{n}}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}.$ (4.16)
@@
$\quad+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-1}\left(1+\left\|{\Phi^{n}}\right\|^{4}\right)\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right]+\frac{1}{3}\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t\right].$ (4.17)
@@
$\leq C\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-1}\Big(1+\left\|{\Phi^{n}}\right\|^{4}\Big)\Big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right]$
$\quad+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\Big(1+\left\|{\Phi^{m}}\right\|^{4}+\left\|{\Phi^{n}}\right\|^{4}+\left|{\mathcal{A}\Phi^{m}}\right|^{2}+\lambda_{n}^{-\frac{1}{2}}\left\|{\Phi^{n}}\right\|^{2}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right]$
$\leq C\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\Big(1+\left|{\mathcal{A}\Phi^{m}}\right|^{2}\Big)\mathrm{d}t\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-\frac{1}{2}}\Big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right],$
@@
We are now in the position to apply the Gronwall lemma for stochastic processes (Lemma A.4) with $u=\left\|{\Phi^{m,n}}\right\|^{2}$, $v=\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$, $w=\lambda_{n}^{-\frac{1}{2}}\big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\big)$, and $r=1+\left|{\mathcal{A}\Phi^{m}}\right|^{2}$. This results in
$\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi^{m,n}}\right\|^{2}\right]+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t\right]\leq C\mathbb{E}\left[\left\|{\Phi^{0}}\right\|^{2}\right]+C\lambda_{n}^{-\frac{1}{2}},$
@@
$\mathds{1}_{\Omega_{\ell}}\int_{0}^{t\wedge\tau}\left\langle P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}}),v\right\rangle\mathrm{d}s$ $\rightharpoonup\int_{0}^{t\wedge\tau}\left\langle B(\Phi),v\right\rangle\mathrm{d}s,$
@@ 5.1. Maximal strong solution for $d\leq 3$
The existence of a maximal strong pathwise solution in the sense of Definition 4.3 will be shown in this section. The argument follows closely the idea in [8, Section 5.3.4]
@@
There exists a unique maximal strong pathwise solution $(\Phi,\xi)$ to (1.1) and a sequence of stopping times $\{\tau_{n}\}_{n\in\mathbb{N}}$ announcing $\xi$. Moreover, $(\Phi,\xi)$ is a local weak pathwise solution.
@@ 5.2. Global strong solution for $d=2$
Acknowledgements
The authors acknowledge financial support through the Australian Research Council’s Discovery Projects funding scheme (project DP200101866). Agus L. Soenjaya is supported by the Australian Government Research Training Program (RTP) Scholarship awarded at the University of New South Wales, Sydney.
The authors would like to thank Ben Goldys for valuable discussions and suggestions related to this work.
@@ Appendix A Auxiliary results
    Then there exists a stopping time $\tau$ with $\mathbb{P}(0<\tau\leq T)=1$ and a process $v$ with $v(\,\cdot\,\wedge\tau)\in\mathcal{E}(\tau)$ such that
    $\left|{v_{n_{l}}-v}\right|_{\mathcal{E}(\tau)}\to 0,\;\text{ a.s. for some subsequence $n_{l}\uparrow\infty$,}$
@@
Suppose that $X$ is a separable Banach space and let $D\subset X$ be a dense subset. Let $X^{\prime}$ be the dual of $X$ and denote the dual pairing between $u\in X$ and $v\in X^{\prime}$ by $\left\langle u,v\right\rangle$. Assume that $(E,\mathcal{E},\mu)$ is a finite measure space. Let $p\in(1,\infty)$. Suppose that $\{v_{n}\}_{n\in\mathbb{N}}$ is a uniformly bounded sequence in $L^{p}(E,X^{\prime})$ and
$\left\langle v_{n},y\right\rangle\to\left\langle v,y\right\rangle,\;\text{ $\mu$-a.e. for all $y\in D$.}$
Then $v_{n}\overset{\ast}{\rightharpoonup}v$ in $L^{p}(E,X^{\prime})$.
@@
Let $\alpha_{1}$ and $\alpha_{2}$ be real numbers such that $\alpha_{1}<\alpha_{2}$. Let $\mathcal{H}_{n}:=\text{span}\{e_{1},e_{2},\ldots,e_{n}\}$, where $\{e_{k}\}_{k\in\mathbb{N}}$ is an orthonormal basis for $\mathcal{H}$ consisting of eigenfunctions of $\mathcal{A}$. Let $P_{n}:\mathcal{H}\to\mathcal{H}_{n}$ be the projection operator and let $Q_{n}=I-P_{n}$, where $I$ is the identity operator. Then for any $v\in\mathrm{D}(\mathcal{A}^{\alpha_{2}})$,
@@
Let $T>0$ be fixed. Suppose that $u,v,w,r:[0,T)\times\Omega\to\mathbb{R}$ are real-valued, non-negative stochastic processes. Let $\tau<T$ be a stopping time such that
$\mathbb{E}\left[\int_{0}^{\tau}(r(t)u(t)+w(t))\,\mathrm{d}t\right]<\infty.$
@@
$\int_{0}^{\tau}r(t)\,\mathrm{d}t\leq\kappa,\;\text{ a.s.}$
@@
$\mathbb{E}\left[\sup_{s\in[\tau_{a},\tau_{b}]}u(s)+\int_{\tau_{a}}^{\tau_{b}}v(t)\,\mathrm{d}t\right]\leq C_{0}\mathbb{E}\left[u(\tau_{a})+\int_{\tau_{a}}^{\tau_{b}}(r(t)u(t)+w(t))\,\mathrm{d}t\right].$

### v2

@@ 3.3. Stochastic Boussinesq model for the Bénard convection
Since $\mathcal{S}$ has the same form as (3.4), it remains to verify (2.7), (2.8), (2.9), and (2.10) for the map $\mathcal{F}$. Firstly, by Hölder’s inequality and Sobolev embedding,
$\left|{\mathcal{F}(\Phi)}\right|^{2}$ $\leq C\left|{\theta}\right|^{2}+C\left|{u}\right|_{L^{4}}^{2}\left|{\nabla\phi}\right|_{L^{4}}^{2}\leq C\left\|{\Phi}\right\|^{2},$
which implies (2.7). Next, by Hölder’s inequality and the Sobolev embedding,
@@
which yields (2.8). Inequality (2.9) can be shown similarly. Finally, by interpolation and Gagliardo–Nirenberg inequalities,
@@
$\leq C\left\|{\Phi}\right\|\left|{\mathcal{A}\Phi}\right|,$
which implies (2.10), thus completing the verification of condition (R).
@@ 3.5. Stochastic micropolar and magneto-micropolar fluid systems
The existence of strong solutions to the deterministic magneto-micropolar fluid system is shown in [40, 49]. In the stochastic case, the existence of martingale weak solutions to the system is proven in [50, 51]. This paper establishes the existence and uniqueness of local pathwise strong solutions to the stochastic magneto-micropolar fluid system for $d\leq 3$. This pathwise strong solution is shown to be global when $d\leq 2$.
@@ 4. Existence and uniqueness of a local strong solution
The pair $(\Phi,\tau)$ is a local weak pathwise solution of (1.1) with initial data $\Phi_{0}$ if $\tau$ is a strictly positive stopping time and $\Phi(\,\cdot\wedge\tau)$ is a predictable process in $\mathcal{V}^{\prime}$ such that
@@
$\Phi(t\wedge\tau)+\int_{0}^{t\wedge\tau}\big(\mathcal{A}\Phi+\mathcal{B}(\Phi)+\mathcal{R}(\Phi)\big)\,\mathrm{d}s=\Phi_{0}+\int_{0}^{t\wedge\tau}g(\Phi)\,\mathrm{d}W\quad\text{in $\mathcal{V}^{\prime}$}.$ (4.2)
@@
A local weak pathwise solution $(\Phi,\tau)$ is said to be a local strong pathwise solution of (1.1) with initial data $\Phi_{0}$ if $\Phi(\,\cdot\wedge\tau)$ is a predictable process in $\mathcal{H}$ such that
@@
and that for any $t>0$, the equality (4.2) holds in $\mathcal{H}$.
@@
$\leq C\lambda_{n}^{-\frac{1}{4}}\left|{Q_{n}\mathcal{B}(\Phi^{n})}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-\frac{1}{4}}\left\|{\Phi^{n}}\right\|^{2}\left|{\mathcal{A}\Phi^{n}}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}.$ (4.13)
@@
$\leq C\lambda_{n}^{-\frac{1}{4}}\left|{Q_{n}\mathcal{R}(\Phi^{n})}\right|_{\mathrm{D}(\mathcal{A}^{\frac{1}{8}})}^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$
$\leq C\lambda_{n}^{-\frac{1}{4}}\left(1+\left\|{\Phi^{n}}\right\|^{4}\right)\left|{\mathcal{A}\Phi^{n}}\right|^{2}+\frac{1}{9}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}.$ (4.16)
@@
$\quad+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-\frac{1}{4}}\left(1+\left\|{\Phi^{n}}\right\|^{4}\right)\left|{\mathcal{A}\Phi^{n}}\right|^{2}\mathrm{d}t\right]+\frac{1}{3}\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t\right].$ (4.17)
@@
$\leq C\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-\frac{1}{4}}\Big(1+\left\|{\Phi^{n}}\right\|^{4}\Big)\Big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right]$
$\quad+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\Big(1+\left\|{\Phi^{m}}\right\|^{4}+\left\|{\Phi^{n}}\right\|^{4}+\left|{\mathcal{A}\Phi^{m}}\right|^{2}+\lambda_{n}^{-\frac{1}{4}}\left\|{\Phi^{n}}\right\|^{2}\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right]$
$\leq C\mathbb{E}\left[\left\|{\Phi^{m,n}(\tau_{a})}\right\|^{2}\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left\|{\Phi^{m,n}}\right\|^{2}\Big(1+\left|{\mathcal{A}\Phi^{m}}\right|^{2}\Big)\mathrm{d}t\right]+C\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\lambda_{n}^{-\frac{1}{4}}\Big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\Big)\mathrm{d}t\right],$
@@
We are now in the position to apply the Gronwall lemma for stochastic processes (Lemma A.4) with $b=1+\left|{\mathcal{A}\Phi^{m}}\right|^{2}$, $u=\left\|{\Phi^{m,n}}\right\|^{2}$, $v=\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}$, and $w=\lambda_{n}^{-\frac{1}{4}}\big(1+\left|{\mathcal{A}\Phi^{n}}\right|^{2}\big)$. This results in
$\mathbb{E}\left[\sup_{t\in[\tau_{a},\tau_{b}]}\left\|{\Phi^{m,n}}\right\|^{2}\right]+\mathbb{E}\left[\int_{\tau_{a}}^{\tau_{b}}\left|{\mathcal{A}\Phi^{m,n}}\right|^{2}\mathrm{d}t\right]\leq C\mathbb{E}\left[\left\|{\Phi^{0}}\right\|^{2}\right]+C\lambda_{n}^{-\frac{1}{4}},$
@@
$\mathds{1}_{\Omega_{\ell}}\int_{0}^{t\wedge\tau}\left\langle P_{n_{\ell}}\mathcal{B}(\Phi^{n_{\ell}}),v\right\rangle\mathrm{d}s$ $\rightharpoonup\int_{0}^{t\wedge\tau}\left\langle\mathcal{B}(\Phi),v\right\rangle\mathrm{d}s,$
@@ 5.1. Maximal strong solution for $d\leq 3$
The existence of a maximal strong pathwise solution in the sense of Definition 4.3 will be shown in this section. The argument follows closely the idea in [7, Section 5.3.4]
@@
There exists a unique maximal strong pathwise solution $(\Phi,\xi)$ to (1.1) and a sequence of stopping times $\{\tau_{n}\}_{n\in\mathbb{N}}$ announcing $\xi$.
@@ Appendix A Auxiliary results
    Then there exists a stopping time $\tau$ with $\mathbb{P}(0<\tau\leq T)=1$, a process $v$ with $v(\,\cdot\,\wedge\tau)\in\mathcal{E}(\tau)$, and a subsequence $\{v_{n_{l}}\}$ such that
    $\left|{v_{n_{l}}-v}\right|_{\mathcal{E}(\tau)}\to 0,\;\text{ a.s. when $l\to\infty$,}$
@@
Suppose that $X$ is a separable Banach space and let $Y\subset X$ be a dense subset. Let $X^{\prime}$ be the dual of $X$ and denote the dual pairing between $u\in X$ and $v\in X^{\prime}$ by $\left\langle u,v\right\rangle$. Assume that $(E,\mathcal{E},\mu)$ is a finite measure space. Let $p\in(1,\infty)$. Suppose that $\{v_{n}\}_{n\in\mathbb{N}}$ is a bounded sequence in $L^{p}(E,X^{\prime})$ and
$\left\langle v_{n},y\right\rangle\to\left\langle v,y\right\rangle,\;\text{ $\mu$-a.e. for all $y\in Y$.}$ (A.7)
Then $v_{n}\overset{\ast}{\rightharpoonup}v$ in $L^{p}(E,X^{\prime})$, i.e., (A.7) holds for all $y\in X.$
@@
Let $\mathcal{H}$ and $\mathcal{A}$ be defined as in Subsection 2.2. Let $\mathcal{H}_{n}:=\text{span}\{e_{1},e_{2},\ldots,e_{n}\}$, where $\{e_{k}\}_{k\in\mathbb{N}}$ is an orthonormal basis for $\mathcal{H}$ consisting of eigenfunctions of $\mathcal{A}$. Let $P_{n}:\mathcal{H}\to\mathcal{H}_{n}$ be the projection operator and let $Q_{n}=I-P_{n}$, where $I$ is the identity operator. Then for any real numbers $\alpha_{1}$ and $\alpha_{2}$, $\alpha_{1}<\alpha_{2}$, if $v\in\mathrm{D}(\mathcal{A}^{\alpha_{2}})$, then
@@
Let $T>0$ be fixed. Suppose that $b,u,v,w:[0,T)\times\Omega\to\mathbb{R}$ are real-valued, non-negative stochastic processes. Let $\tau<T$ be a stopping time such that
$\mathbb{E}\left[\int_{0}^{\tau}(b(t)u(t)+w(t))\,\mathrm{d}t\right]<\infty.$
@@
$\int_{0}^{\tau}b(t)\,\mathrm{d}t\leq\kappa,\;\text{ a.s.}$
@@
$\mathbb{E}\left[\sup_{s\in[\tau_{a},\tau_{b}]}u(s)+\int_{\tau_{a}}^{\tau_{b}}v(t)\,\mathrm{d}t\right]\leq C_{0}\mathbb{E}\left[u(\tau_{a})+\int_{\tau_{a}}^{\tau_{b}}(b(t)u(t)+w(t))\,\mathrm{d}t\right].$
@@
Statements and declarations
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Funding statement
The authors acknowledge financial support through the Australian Research Council’s Discovery Projects funding scheme (projects DP200101866 and DP220101811).
Agus L. Soenjaya is supported by the Australian Government Research Training Program (RTP) Scholarship awarded at the University of New South Wales, Sydney.
Data availability
No data are associated with this article.